If the base of a triangle is 8x+16 yards, and its height is 1 over x+2 yards (written as a fraction), then what is the area of the triangle?
Thank you!

Answer by Theo(13342)

b = 8x+16
h = 1/(x+2)
a = (1/2)*b*h

your formula is:

a = (1/2)*b*h = (1/2)*(8x+16)*(1/(x+2)

(1/2)*(8x+16) = (4x+8)

(4x+8) = 4*(x+2)

your equation becomes:

(4*(x+2))*(1/(x+2)

this equals 4 because (x+2)/(x+2) = 1

your area is equal to 4 regardless of the value of x.

when x = 1, your formula of

a = (1/2)*(8x+16)*(1/(x+2) becomes:

a = (1/2)*24*(1/3) = 12/3 = 4

when x = 4, your formula of

a = (1/2)*(8x+16)*(1/(x+2) becomes:

a = (1/2)*(48)*(1/6) = 24/6 = 4

etc.

your answer is the area = 4 yards